OpenStudy (anonymous):

2. William said that when you multiply three positive and negative decimals, the answer changes depending on where you place the parentheses. Use these numbers to answer the questions: –2.4, –0.63, 1.8, –3.42 (a) Is William’s statement always true, sometimes true, or always false? Write two pairs of equations to support your answer. (b) Write a rule describing the results when one multiplies three positive and negative decimals and changes the order of multiplication using parentheses. please help with my two questions =^.^=

OpenStudy (anonymous):

@Michele_Laino

OpenStudy (michele_laino):

hint: the multiplication of rational numbers has the commutative property, namely: \[\huge \begin{gathered} \left( {a \cdot b} \right) \cdot c = a \cdot \left( {b \cdot c} \right) = \left( {c \cdot a} \right) \cdot b \hfill \\ \hfill \\ a,b,c \in \mathbb{Q} \hfill \\ \end{gathered} \]